a 38 N horizontal force pushes on a 3.7 Kg mass resting on a horizonal surface. The surface exerts a friction of 22 N against the motion. The acceleration of the 3.7 kg Mass is
Answer #1

According to Newton's Second law of motion:

Fnet =ma

Since friction is an opposing force we have Fnet =38 - 22 = 16 N

Given mass = 3.7 kg, then acceleration is given by:

Fnet 16 2 4.32m/s 3.7 a = m